# Disentangling Latent Shifts of In-Context Learning Through Self-Training

### v1

@@ 3.2 Adapter Arithmetic
To overcome the limitations of context window sizes and efficiently handle extensive demonstration sets in ICL, we employ adapter arithmetic within sticl. sticl achieves this by fine-tuning separate adapters for each demonstration subset, with each adapter encoding the latent shift corresponding to its subset. These adapters are then merged by summing their parameters (Chitale et al., 2023), resulting in a single adapter that integrates knowledge from all subsets. Partitioning demonstrations into smaller subsets allows for better use of long contexts and effectively extending them without exceeding window limits or altering the base LLM architecture. Additionally, distributing the prompt across multiple adapters optimizes GPU utilization, fitting the entire prompt on a single GPU during inference and reducing memory constraints.
@@
Table 4 shows the ID generalization scores of ICV, Batch-ICL, and sticl in fusing knowledge from multiple demonstration subsets, specifically using $2$, $4$, and $8$ subsets of $16$ demonstrations each. sticl-S consistently outperforms baseline methods, demonstrating its ability to fuse knowledge from different subsets. This success parallels broader trends in knowledge fusion within LLMs Wan et al. (2024). Moreover, this form of adapter arithmetic aligns with recent advances in task arithmetic, where merging task-specific parameters promotes generalization across multiple tasks (Ilharco et al., 2023; Ortiz-Jimenez et al., 2023). In our case, this approach effectively improves generalization and stability when fusing demonstration subsets within the same task.
@@
Table 4: ID generalization scores of knowledge fusion for Llama 3 (8B). The scores are averaged over $10$ runs with standard deviations shown as subscripts. The table compares the effectiveness of knowledge fusion from $2$, $4$, and $8$ subsets of $16$ demonstrations. The highest scores are in bold.
@@
GLUE MMLU Demonstrations Method rte sst qnli mnli cola mrpc qqp math misc $\mathbf{2\times 16}$ ICV $75.2_{4.3}$ $93.6_{1.9}$ $77.6_{5.9}$ $69.2_{3.7}$ $58.3_{3.5}$ $74.2_{2.4}$ $70.6_{2.7}$ $45.5_{3.7}$ $72.5_{2.9}$ Batch-ICL $80.2_{3.6}$ $95.3_{1.8}$ $80.2_{5.8}$ $72.3_{3.0}$ $61.2_{3.1}$ $76.3_{2.0}$ $72.6_{2.4}$ $43.5_{2.9}$ $83.0_{3.6}$ sticl-S $\textbf{87.1}_{1.6}$ $\textbf{96.4}_{1.3}$ $\textbf{81.5}_{5.0}$ $\textbf{75.5}_{2.5}$ $\textbf{68.4}_{1.8}$ $\textbf{78.5}_{1.4}$ $\textbf{74.1}_{1.6}$ $\textbf{51.5}_{1.6}$ $\textbf{89.5}_{2.0}$ $\mathbf{4\times 16}$ ICV $78.3_{3.6}$ $94.6_{1.8}$ $79.3_{5.5}$ $71.2_{3.1}$ $60.3_{3.3}$ $75.6_{2.2}$ $72.3_{2.4}$ $47.5_{3.5}$ $76.5_{3.8}$ Batch-ICL $84.4_{3.3}$ $96.4_{1.5}$ $82.4_{5.2}$ $74.3_{2.5}$ $64.2_{2.8}$ $78.3_{1.6}$ $74.3_{2.1}$ $45.5_{2.6}$ $84.5_{3.3}$ sticl-S $\textbf{88.4}_{2.3}$ $\textbf{97.5}_{0.7}$ $\textbf{83.6}_{4.4}$ $\textbf{77.3}_{2.2}$ $\textbf{71.4}_{1.5}$ $\textbf{79.6}_{0.7}$ $\textbf{75.2}_{1.3}$ $\textbf{53.5}_{1.4}$ $\textbf{91.0}_{1.7}$ $\mathbf{8\times 16}$ ICV $81.3_{2.8}$ $95.6_{1.5}$ $81.8_{5.0}$ $73.3_{2.7}$ $61.3_{2.4}$ $77.3_{1.7}$ $73.8_{2.0}$ $47.5_{2.9}$ $78.0_{3.5}$ Batch-ICL $85.6_{2.5}$ $96.7_{1.1}$ $83.8_{4.5}$ $75.8_{2.1}$ $65.3_{2.1}$ $79.8_{1.3}$ $75.8_{1.8}$ $45.5_{2.0}$ $84.0_{2.5}$ sticl-S $\textbf{92.8}_{0.8}$ $\textbf{98.1}_{0.2}$ $\textbf{87.9}_{2.5}$ $\textbf{81.3}_{0.9}$ $\textbf{74.1}_{0.6}$ $\textbf{82.8}_{0.4}$ $\textbf{78.9}_{0.5}$ $\textbf{57.0}_{0.5}$ $\textbf{93.0}_{0.7}$
@@ 4 Analysis of Weak-to-Strong Generalization
(a) Histogram of approximated Lipschitz constants
(b) Rate of pseudo-label correction over epochs
(c) Rate of corrected ID examples
(d) Rate of corrected OOD examples
@@
Figure 2: Empirical analysis of sticl-S on the aggregated GLUE datasets for LLaMA-3 (8B): (a) Histogram of approximated Lipschitz constants across datasets, computed as the Frobeinus norm of the input-output Jacobian matrix; (b) Rate of pseudo-label correction over training epochs with examples from the unlabeled dataset used for self-training. Shaded areas indicate the standard deviation over $10$ runs; (c) and (d) Corrected and corrupted prediction rates for (c) ID examples and (d) OOD examples, based on the Euclidean distance to the closest correctly pseudo-labeled neighbor (normalized to $[0,1]$). There are $10$ bins ranging from the interval of $[0,0.1]$ to $[0.9,1]$. Error bars denote the standard deviation over $10$ runs.
@@
Building on the observation that sticl consistently outperforms its teacher, standard ICL, we hypothesize that weak-to-strong generalization may be driving these improvements, where the model’s ability to generalize strengthens progressively from weaker signals. To explore this further, we conduct an empirical analysis of sticl-S with Llama 3 (8B) on aggregated examples from all GLUE datasets, treating them as a single, unified dataset.
4.1 Local Consistency
@@
A crucial prerequisite for successful weak-to-strong generalization is the student’s ability to maintain stable outputs under small perturbations of the input, i.e., robustness to input variations. A low Lipschitz constant serves as a key indicator of this stability, as it bounds the maximum change in the model’s output for any change in its input (Khromov & Singh, 2024). However, calculating the exact Lipschitz constant for LLMs is intractable. To approximate it, we leverage the relationship between the Lipschitz constant and the input-output Jacobian matrix of a neural network. Specifically, we compute the Frobenius norm of the Jacobian matrix as a tractable proxy, given its relationship to the spectral norm, which is a known lower bound for the Lipschitz constant (Dherin et al., 2022) (cf. Appendix B for theoretical details). Figure 2(a) presents the distribution of the approximated Lipschitz constants (normalized to $[0,1]$) for sticl, PBFT, and ICL, providing a proxy for local consistency. sticl exhibits a notably lower Lipschitz constant than PBFT and ICL, underscoring its local consistency.
4.2 Pseudo-Label Correction and Coverage Expansion
@@
Pseudo-label correction, where the student model revises the labels predicted by the teacher model, is a fundamental mechanism that drives weak-to-strong generalization (Lang et al., 2024). This process is closely tied to the model’s ability to establish local consistency within the representation space, where accurate predictions in confident regions propagate corrections to neighboring, less certain areas, fostering local-to-global consistency throughout training. Figure 2(b) shows how the rate of corrected pseudo-labels evolves during training on GLUE datasets. As training progresses, the percentage of corrected pseudo-labels steadily increases, showcasing sticl’s capacity to exhibit weak-to-strong generalization. Notably, the rate of pseudo-label correction plateaus faster for simpler datasets like sst and qnli, which have lower linguistic variability.
@@
The mechanism of pseudo-label correction ties into the phenomenon of coverage expansion – where the model generalizes beyond the regions covered by pseudo-labels Lang et al. (2024). We hypothesize that the core of sticl’s ability to generalize effectively is anchored in coverage expansion, which enables local corrections to propagate globally, creating a ripple effect across the representation space. To understand this dynamic, we analyze which unseen evaluation points are corrected by clustering them based on their proximity to the nearest correctly pseudo-labeled neighbor in $\mathcal{D}_{\text{unlab}}$. This is quantified by computing the Euclidean distance between the model’s representations at the final hidden states, with evaluation points categorized into ten bins based on their normalized distance from the correct neighbor, spanning the range $[0,1]$. Figure 2(c) illustrates the rate of prediction flips within these bins, where a flip refers to either correcting an incorrect prediction or corrupting a correct one. The rate of corrected predictions shows a strong negative correlation with the distance to the nearest correctly labeled neighbor, as indicated by a Pearson correlation coefficient of $-0.968$, while corrupted predictions are more frequent in regions lacking nearby correct pseudo-labels.
Coverage expansion shows its effects even on OOD data. Figure 2(d), the counterpart to Figure 2(c), shows the rate of flipped predictions for OOD data. Although the impact is reduced, a similar correction pattern persists, with a Pearson correlation of $-0.916$. This consistency across domains highlights the model’s ability to propagate accurate predictions not only within the training domain but also across OOD data.
@@ 5 Related Work
ICL theory.
The understanding of ICL has shifted from a traditional task-learning framework to one focused on task identification. Wies et al. (2023) argue that ICL operates by recognizing latent tasks embedded within a model’s pre-training, allowing for efficient performance on new tasks. Building on this, Hoogland et al. (2024) suggest that ICL in transformers progresses through distinct developmental stages, offering deeper insights into how models adapt to unfamiliar contexts. Li et al. (2023) further empirically show that ICL predictions become more resilient to input perturbations with longer prompts and that training on noisy data enhances stability. Despite these theoretical breakthroughs, ICL remains vulnerable to the selection and ordering of demonstrations (Li et al., 2024; Lu et al., 2021). Moreover, Kossen et al. (2024) highlight ICL’s biases rooted in pre-training data, revealing that models do not always uniformly leverage in-context information.
Disentaglement of latent shifts.
Research into the inner workings of ICL has revealed how transformers process demonstrations to form task representations. Hendel et al. (2023) and Liu et al. (2023) show that transformers can compress demonstration examples into a task vector, which efficiently directs the model to generate context-appropriate outputs for queries. These task vectors are created during a forward pass, capturing the latent shift induced by the demonstrations. Building on this, Dai et al. (2023) explore using linear attention to compute virtual gradients, simulating the effect of gradient-based learning within the model. Similarly, Todd et al. (2024) use causal mediation analysis to highlight the role of specific attention heads in forming robust task representations in ICL, termed function vectors.
Self-training and weak-to-strong generalization.
Wei et al. (2021) provide a theoretical foundation for self-training, showing that under the assumption of coverage expansion, the minimizers of population objectives based on self-training and local consistency regularization achieve high accuracy. Lang et al. (2024) further develop the principle of pseudo-label correction, which occurs when the student model demonstrates strong local consistency. Several works have extended these ideas in the context of LLMs. For instance, Huang et al. (2023) demonstrate that LLMs can enhance their reasoning abilities through self-training without the need for labeled data by generating high-confidence, rationale-augmented answers, which are then used for fine-tuning, leading to improved performance across various tasks. In the same vein, Qu et al. (2024) propose recursive introspection for self-improvement, and Wang et al. (2024) introduce self-taught evaluators, showing how LLMs can autonomously refine and improve their outputs over time.
@@ 6 Conclusion
We tackled the challenges of stability and long-context handling that arise when processing multiple demonstrations in ICL within LLMs. To address these issues, we introduced sticl (Self-Training ICL), a method that disentangles the latent shifts induced by demonstrations from those of the query, leveraging a teacher-student framework. sticl encodes these latent shifts into an adapter module, enabling the student model to handle queries without requiring demonstrations in the input. Moreover, sticl allows efficient handling of large demonstration sets by chunking them into manageable subsets, each processed through separate adapter modules. This not only reduces the instability caused by demonstration selection and ordering but also alleviates the context window limitations inherent in transformer-based models. We demonstrated that sticl exhibits weak-to-strong generalization by refining pseudo-labels through progressive corrections, expanding from local consistency to a more comprehensive coverage across the representation space. Our empirical evaluation of sticl showed that it consistently outperforms traditional ICL methods, significantly improving generalization and stability across diverse datasets. These findings underscore the effectiveness of self-training as a promising strategy for improving ICL performance.
@@ Appendix A Dual Form of ICL
We offer a detailed derivation of (5), originally introduced by Dai et al. (2023), expanding on the key intermediate steps for clarity, which were not explicitly covered in the original work. The goal is to decompose the attention head output into separate components corresponding to the demonstrations and the query, thereby disentangling the latent shifts induced by ICL.
@@ A.1 Starting Point
  • $\mathbf{X}_{q}\in\mathbb{R}^{d_{\text{model}}\times N_{q}}$ is the matrix of previous query token representations up to time $t-1$;
@@
  • $[\mathbf{X}_{d};\mathbf{X}_{q}]$ is the concatenation of $\mathbf{X}_{d}$ and $\mathbf{X}_{q}$ along the sequence dimension.

### v2

@@ 3.3 Adapter Arithmetic
To overcome the limitations of context window sizes and efficiently handle extensive demonstration sets in ICL, we employ adapter arithmetic within wilda. This is achieved by fine-tuning separate adapters for each demonstration subset, with each adapter encoding the latent shift corresponding to its subset. Following the approach of Chitale et al. [6], these adapters are merged by summing their parameters, producing a single adapter that integrates knowledge from all subsets. Partitioning demonstrations into smaller subsets enables more effective use of the available context window, allowing models to incorporate more demonstrations without exceeding length limits or modifying the base LLM architecture. Additionally, distributing the prompt across multiple adapters improves GPU utilization by fitting it on a single GPU and reducing memory overhead during inference.
@@
Table 4 shows the ID generalization scores of ICV, Batch-ICL, and wilda in fusing knowledge from multiple demonstration subsets, specifically using $2$, 44, and $8$ subsets of $16$ demonstrations each. wilda-S consistently outperforms baseline methods, highlighting its effectiveness in knowledge fusion across subsets [35]. Moreover, this form of adapter arithmetic aligns with recent advances in task arithmetic, where merging task-specific parameters promotes generalization across multiple tasks [18, 30]. In our case, this approach effectively improves generalization and stability when fusing demonstration subsets within the same task.
@@
Table 4: ID generalization scores of knowledge fusion for Llama 3. The scores are averaged over $10$ runs with standard deviations shown as subscripts. The table compares the effectiveness of knowledge fusion from 22, 44, and $8$ subsets of $16$ demonstrations. The highest scores are in bold.
@@
GLUE MMLU Demonstrations Method rte sst qnli mnli cola mrpc qqp math misc $\mathbf{2\times 16}$ ICV $75.2_{4.3}$ $93.6_{1.9}$ $77.6_{5.9}$ $69.2_{3.7}$ $58.3_{3.5}$ $74.2_{2.4}$ $70.6_{2.7}$ $45.5_{3.7}$ $72.5_{2.9}$ Batch-ICL $80.2_{3.6}$ $95.3_{1.8}$ $80.2_{5.8}$ $72.3_{3.0}$ $61.2_{3.1}$ $76.3_{2.0}$ $72.6_{2.4}$ $43.5_{2.9}$ $83.0_{3.6}$ wilda-S $\textbf{87.1}_{1.6}$ $\textbf{96.4}_{1.3}$ $\textbf{81.5}_{5.0}$ $\textbf{75.5}_{2.5}$ $\textbf{68.4}_{1.8}$ $\textbf{78.5}_{1.4}$ $\textbf{74.1}_{1.6}$ $\textbf{51.5}_{1.6}$ $\textbf{89.5}_{2.0}$ $\mathbf{4\times 16}$ ICV $78.3_{3.6}$ $94.6_{1.8}$ $79.3_{5.5}$ $71.2_{3.1}$ $60.3_{3.3}$ $75.6_{2.2}$ $72.3_{2.4}$ $47.5_{3.5}$ $76.5_{3.8}$ Batch-ICL $84.4_{3.3}$ $96.4_{1.5}$ $82.4_{5.2}$ $74.3_{2.5}$ $64.2_{2.8}$ $78.3_{1.6}$ $74.3_{2.1}$ $45.5_{2.6}$ $84.5_{3.3}$ wilda-S $\textbf{88.4}_{2.3}$ $\textbf{97.5}_{0.7}$ $\textbf{83.6}_{4.4}$ $\textbf{77.3}_{2.2}$ $\textbf{71.4}_{1.5}$ $\textbf{79.6}_{0.7}$ $\textbf{75.2}_{1.3}$ $\textbf{53.5}_{1.4}$ $\textbf{91.0}_{1.7}$ $\mathbf{8\times 16}$ ICV $81.3_{2.8}$ $95.6_{1.5}$ $81.8_{5.0}$ $73.3_{2.7}$ $61.3_{2.4}$ $77.3_{1.7}$ $73.8_{2.0}$ $47.5_{2.9}$ $78.0_{3.5}$ Batch-ICL $85.6_{2.5}$ $96.7_{1.1}$ $83.8_{4.5}$ $75.8_{2.1}$ $65.3_{2.1}$ $79.8_{1.3}$ $75.8_{1.8}$ $45.5_{2.0}$ $84.0_{2.5}$ wilda-S $\textbf{92.8}_{0.8}$ $\textbf{98.1}_{0.2}$ $\textbf{87.9}_{2.5}$ $\textbf{81.3}_{0.9}$ $\textbf{74.1}_{0.6}$ $\textbf{82.8}_{0.4}$ $\textbf{78.9}_{0.5}$ $\textbf{57.0}_{0.5}$ $\textbf{93.0}_{0.7}$
@@ 4 Analysis of Weak-to-Strong Generalization
Figure 2: Empirical analysis of wilda-S on the aggregated GLUE datasets for Llama 3: (a) Histogram of approximated Lipschitz constants across datasets, computed as the Frobenius norm of the input–output Jacobian matrix; (b) Rate of pseudo-label correction over training epochs (shaded areas indicate the standard deviation over $10$ runs); Corrected and corrupted prediction rates for (c) ID examples and (d) OOD examples, based on the normalized Euclidean distance to the nearest correctly pseudo-labeled neighbor. Error bars indicate standard deviation over $10$ runs.
@@
Building on the observation that wilda consistently outperforms its teacher, standard ICL, we hypothesize that W2S generalization may be driving these improvements, where the model’s ability to generalize strengthens progressively from weaker signals. To explore this further, we conduct an empirical analysis of wilda-S with Llama 3 on aggregated examples from all GLUE datasets, treating them as a single, unified dataset. We focus on the wilda-S variant due to its consistently strong performance and stability across prior experiments.
@@
A crucial prerequisite for successful W2S generalization is the student’s ability to maintain stable outputs under small perturbations of the input, i.e., robustness to input variations. A low Lipschitz constant serves as a key indicator of this stability, as it bounds the maximum change in the model output for any change in its input [20]. However, calculating the exact Lipschitz constant for LLMs is intractable. To approximate it, we leverage the relationship between the Lipschitz constant and the input–output Jacobian matrix of a neural network. Specifically, we compute the Frobenius norm of the Jacobian matrix as a tractable proxy, given its relationship to the spectral norm, which is a known lower bound for the Lipschitz constant [8] (see Appendix B). Figure 2(a) shows a histogram of the approximated Lipschitz constants (normalized to $[0,1]$) for wilda, PBFT, and ICL. wilda exhibits a notably lower Lipschitz constant than PBFT and ICL, which reflects its stronger local consistency.
@@
Pseudo-label correction, where the student model revises the labels predicted by the teacher, builds directly on local consistency and plays a central role in enabling W2S generalization [22]. When the model exhibits stable behavior under small input perturbations, as captured by a low Lipschitz constant, it is more likely to propagate corrections reliably across neighboring inputs in representation space. This local consistency forms the foundation for accurate correction of noisy pseudo-labels in high-confidence regions. As corrected labels accumulate, they create a foundation for generalization to nearby, low-confidence examples – gradually expanding the model’s coverage and facilitating a transition from local consistency to broader generalization. Figure 2(b) shows how the rate of corrected pseudo-labels evolves during training on GLUE datasets. As training progresses, the percentage of corrected pseudo-labels steadily increases, demonstrating wilda’s capacity to exhibit W2S generalization. Notably, the rate of pseudo-label correction plateaus faster for simpler datasets like sst and qnli, which have lower linguistic variability.
@@
The mechanism of pseudo-label correction ties into the phenomenon of coverage expansion, where the model generalizes beyond the regions covered by pseudo-labels [22]. We hypothesize that wilda ’s generalization ability is supported by coverage expansion, where local corrections gradually influence nearby examples in representation space. The model’s coverage expands incrementally through a ripple effect, with high-confidence predictions influencing nearby examples while remaining grounded in regions supported by learned corrections. To understand this dynamic, we analyze which unseen evaluation points are corrected by clustering them based on their proximity to the nearest correctly pseudo-labeled neighbor in $\mathcal{D}_{\text{unlab}}$. This is quantified by computing the Euclidean distance between the model’s representations at the final hidden states, with evaluation points categorized into ten bins based on their normalized distance from the correct neighbor. Figure 2(c) illustrates the rate of prediction flips within these bins, where a flip refers to either correcting an incorrect prediction or corrupting a correct one. The rate of corrected predictions shows a strong negative association with the distance to the nearest correctly labeled neighbor, as measured by the point biserial correlation coefficient of 0.968-0.968. In contrast, corrupted predictions are more frequent in regions lacking nearby correct pseudo-labels. Moreover, coverage expansion also shows its effects on OOD data. Figure 2(d) shows the rate of flipped predictions for OOD data. Although the impact is reduced, a similar correction pattern persists, with a point biserial correlation of 0.916-0.916.
@@ 5 Related Work
Recent perspectives on ICL have shifted from task learning to task identification. Wies et al. [39] argue that ICL works by recognizing latent tasks embedded during pre-training. Hoogland et al. [14] build on this, suggesting that ICL unfolds in developmental stages, shedding light on how models adapt to novel contexts. Li et al. [25] further empirically show that ICL predictions become more resilient to input perturbations with longer prompts and that training on noisy data enhances stability. Despite these theoretical breakthroughs, ICL remains vulnerable to the selection and ordering of demonstrations [24, 29]. Moreover, Kossen et al. [21] highlight ICL’s biases rooted in pre-training data, revealing that models do not always uniformly leverage in-context information.
Research into the inner workings of ICL has revealed how transformers process demonstrations to form task representations. Hendel et al. [12] and Liu et al. [27] show that transformers can compress demonstration examples into a task vector, which efficiently directs the model to generate context-appropriate outputs for queries. These task vectors are created during a forward pass, capturing the latent shift induced by the demonstrations. Building on this, Dai et al. [7] explore using linear attention to compute virtual gradients, simulating the effect of gradient-based learning within the model. Similarly, Todd et al. [33] use causal mediation analysis to highlight the role of specific attention heads in forming robust task representations in ICL, termed function vectors.
Wei et al. [38] provide a theoretical foundation for W2S generalization, showing that, under the assumption of coverage expansion, models optimized for population-level consistency can achieve high accuracy. Lang et al. [22] further advance this view by formalizing the role of pseudo-label correction, which emerges when a model enforces local consistency during training. Building on these principles, several recent works demonstrate how large language models (LLMs) can leverage their own high-confidence outputs to improve performance. For instance, Huang et al. [17] show that rationale-augmented predictions can guide fine-tuning and enhance reasoning abilities without labeled supervision. Similarly, Qu et al. [31] propose recursive introspection for iterative improvement, and Wang et al. [37] introduce self-taught evaluators that enable LLMs to refine their outputs over time.
@@ 6 Conclusion
We tackled the challenges of stability and long-context handling that arise when processing multiple demonstrations in ICL within LLMs. To address these issues, we introduced wilda, a method that disentangles the latent shifts induced by demonstrations from those of the query, leveraging a teacher–student framework. wilda encodes these latent shifts into an adapter module, enabling the student model to handle queries without requiring demonstrations in the input. Moreover, wilda allows efficient handling of large demonstration sets by chunking them into manageable subsets, each processed through separate adapter modules. This not only reduces the instability caused by demonstration selection and ordering but also alleviates the context window limitations inherent in transformer-based models. We demonstrated that wilda exhibits weak-to-strong generalization by refining pseudo-labels through progressive corrections, expanding from local consistency to a more comprehensive coverage across the representation space. Our empirical evaluation of wilda showed that it consistently outperforms traditional ICL methods, significantly improving generalization and stability across diverse natural language understanding datasets.
@@ Appendix A Dual Form of ICL
We offer a detailed derivation of (3), originally introduced by [7], expanding on the key intermediate steps for clarity, which were not explicitly covered in the original work. The goal is to decompose the attention head output into separate components corresponding to the demonstrations and the query, thereby disentangling the latent shifts induced by ICL.
@@ A.1 Starting Point
  • $\mathbf{X}_{q}\in\mathbb{R}^{d_{\text{model}}\times N_{q}}$ is the matrix of previous query token representations up to time t1t-1;
@@
  • $\mathbf{X}_{d};\mathbf{X}_{q}]$ is the concatenation of $\mathbf{X}_{d}$ and $\mathbf{X}_{q}$ along the sequence dimension.